# Modular Control of Biological Networks

### v1

@@ 3 Control via Modularity
Thus,
$\mathcal{D}(F^{\mu})=\mathcal{D}(F_{1}^{\mu_{1}}\rtimes_{P}F_{2}^{\mu_{2}})=\bigsqcup_{\mathcal{C}^{\prime}\in\mathcal{D}(F_{1}^{\mu_{1}})}\mathcal{C}^{\prime}\oplus\mathcal{D}(F^{\mathcal{C}^{\prime},\mu_{2}}_{2})=\mathcal{C}_{1}\oplus\mathcal{D}(F^{\mathcal{C}_{1},\mu_{2}}_{2})=\mathcal{C}_{1}\oplus\mathcal{C}_{2}.$
For the last equality we used the fact that the product of a steady state and a cycle (or vice versa) will result in only one attractor for the combined network. The former is not always true in general because multiplying two attractors (of length greater than 1) might result in several attractors for the composed network due to the attractors starting at different states.
It follows that there is only one attractor of $F^{\mu}$ and that attractor is $\mathcal{C}_{1}\oplus\mathcal{C}_{2}$. Thus, $F$ is stabilized by $\mu=(\mu_{1},\mu_{2})$ and we have $\mathcal{D}(F^{\mu})=\mathcal{C}$.
@@
Given a decomposable network $F=F_{1}\rtimes_{P}F_{2}$, let $\mathcal{C}=\mathcal{C}_{1}\oplus\mathcal{C}_{2}$ be an attractor of $F$, where $\mathcal{C}_{1}\in\mathcal{D}(F_{1})$ and $\mathcal{C}_{2}\in\mathcal{D}(F^{\mathcal{C}_{1}}_{2})$ and at least $\mathcal{C}_{1}$ or $\mathcal{C}_{2}$ is a steady state. If $\mu_{1}$ is a control that stabilizes $F_{1}$ in $\mathcal{C}_{1}$ and $\mu_{2}$ is a control that stabilizes $F^{\mathcal{C}_{1}}_{2}$ in $\mathcal{C}_{2}$, then $\mu=(\mu_{1},\mu_{2})$ is a control that stabilizes $F$ in $\mathcal{C}$.
@@
In Theorem 3.4, we required one of the stabilized attractors to be a steady state in order to be able to combine the controls from the modules. We can remove this requirement from Theorem 3.4 by using the following definition of stabilization for non-autonomous networks, which will guarantee that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ can be combined in a unique way, resulting in a unique attractor of the whole network.
@@
A non-autonomous Boolean network is defined by
@@
Consider a controlled non-autonomous network given by $y(t+1)=\overline{F}_{2}(g(t),y(t),u)$, where $g(t)$ is a trajectory representation of an attractor $\mathcal{C}_{1}$ of an upstream network. We say that a control $\mu_{2}$ stabilizes this network, $F_{2}^{\mathcal{C}_{1}}$ (defined as in Definition 3.7), at an attractor $\mathcal{C}_{2}$ when the resulting network after applying $\mu_{2}$ (denoted here as $F_{2}^{\mathcal{C}_{1},\mu_{2}}$) has $\mathcal{C}_{2}$ as its unique attractor. For non-autonomous networks the definition of unique attractor requires that $\left(g(t),y(t)\right)_{t=0}^{\infty}$ has a unique periodic trajectory up to shifting of $t$ (which is automatically satisfied if $\mathcal{C}_{1}$ or $\mathcal{C}_{2}$ is a steady state).
Example 3.9.
Consider, again, the network $F(x_{1},x_{2},x_{3},x_{4})=(x_{2},x_{1},x_{2}x_{4},x_{3})$, which can be decomposed into $F=F_{1}\rtimes F_{2}$, with $F_{1}(x_{1},x_{2})=(x_{2},x_{1})$ and $F_{2}(x_{3},x_{4})=(x_{4},x_{3})$. Suppose we want to stabilize $F$ in 0110 (which is not an attractor of $F$). Note that the non-autonomous network $\overline{F}_{2}(x_{1},x_{2},x_{3},x_{4})=(x_{2}x_{4},x_{3})$ and $\mathcal{D}(F_{1})=\left\{00,11,\{01,10\}\right\}$.
[…]
  • Finally, the control $\mu=(\mu_{1},\mu_{2})$ stabilizes $F$ at $\mathcal{C}=\mathcal{C}_{1}\oplus\mathcal{C}_{2}=\{0110\}$. Note that $\mathcal{C}$ is a new attractor for $F$.
Theorem 3.4 uses the modular structure of a Boolean network to identify controls that stabilize the network in any desired attractor. In biological applications, the attractors typically correspond to distinct biological phenotypes (defined more rigorously in the next section) and a common question is how to force a network to always transition to only one of these phenotypes. For example, cancer biologists may use an appropriate Boolean network model with the two phenotypes proliferation and apoptosis to identify drug targets (i.e., edge or node controls), which force the system to always undergo apoptosis. The following example illustrates this specific control aspect, described in Corollary 3.5.
Example 3.10.
Consider again the network $F(x_{1},x_{2},x_{3},x_{4})=(x_{2},x_{1},x_{2}x_{4},x_{3})=F_{1}\rtimes F_{2}$ from Example 3.9 with $F_{1}(x_{1},x_{2})=(x_{2},x_{1})$ and $F_{2}(x_{3},x_{4})=(x_{4},x_{3})$. Suppose we want to stabilize $F$ in $1$, which is an attractor of $F$ (but not the only one). Note that the non-autonomous network $\overline{F}_{2}(x_{1},x_{2},x_{3},x_{4})=(x_{2}x_{4},x_{3})$ and $\mathcal{D}(F_{1})=\left\{00,11,\{01,10\}\right\}$. Let $\mathcal{C}_{1}=\{11\}\in\mathcal{D}(F_{1})$.
[…]
Figure 4: Control via modularity and canalization. Once the network is decomposed into modules $F_{1}$, $\cdots$, $F_{n}$, we could override the effect of module $F_{i}$ by the using another module ($F_{k}$ in this case) whose variables are inputs of $f_{x}$ that are located in a higher canalizing hierarchy than the layers containing the variables of $F_{i}$.
@@ 4 Control via Modularity and Canalization
In addition to using the modular structure of the network, we can take advantage of the canalizing structure of the regulatory functions to identify contol targets.
@@
$f(\operatorname{\mathbf{x}})\neq f(\operatorname{\mathbf{x}}\oplus\,e_{i}),$
where $e_{i}$ is the $i$th unit vector. In that case, we also say $f$ depends on $x_{i}$.
@@
[…]
@@
A Boolean function $f(x_{1},\ldots,x_{n})$ is nested canalizing with respect to the permutation $\sigma\in\mathcal{S}_{n}$, inputs $a_{1},\ldots,a_{n}$ and outputs $b_{1},\ldots,b_{n}$, if
$f(x_{1},\ldots,x_{n})=\left\{\begin{array}[c]{ll}b_{1}&x_{\sigma(1)}=a_{1},\\ b_{2}&x_{\sigma(1)}\neq a_{1},x_{\sigma(2)}=a_{2},\\ b_{3}&x_{\sigma(1)}\neq a_{1},x_{\sigma(2)}\neq a_{2},x_{\sigma(3)}=a_{3},\\ \vdots&\vdots\\ b_{n}&x_{\sigma(1)}\neq a_{1},\ldots,x_{\sigma(n-1)}\neq a_{n-1},x_{\sigma(n)}=a_{n},\\ 1\oplus b_{n}&x_{\sigma(1)}\neq a_{1},\ldots,x_{\sigma(n-1)}\neq a_{n-1},x_{\sigma(n)}\neq a_{n}.\end{array}\right.$
@@
where each $M_{i}=\displaystyle\prod_{j=1}^{k_{i}}(x_{i_{j}}+a_{i_{j}})$ is a nonconstant extended monomial, $p_{C}$ is the core polynomial of $f$, and $k=\displaystyle\sum_{i=1}^{r}k_{i}$ is the canalizing depth. Each $x_{i}$ appears in exactly one of $\{M_{1},\ldots,M_{r},p_{C}\}$, and the only restrictions are the following “exceptional cases”:
@@
From Equation 3, we can directly derive an important summary statistic of canalizing functions.
@@
Given a Boolean function $f(x_{1},\ldots,x_{n})$ represented as in Equation  3, we call the extended monomials $M_{i}$ the layers of f and define, as in [29], the layer structure as the vector $(k_{1},\ldots,k_{r})$, which describes the number of variables in each layer. Note that $f$ is nested canalizing if and only if $k_{1}+\cdots+k_{r}=n$.
@@
The Boolean functions $f(x_{1},x_{2},x_{3},x_{4})=x_{1}\wedge(\neg x_{2}\vee(x_{3}\wedge x_{4}))$ and $g(x_{1},x_{2},x_{3},x_{4})=x_{1}\wedge(\neg x_{2}\vee x_{3}\vee x_{4})$ are nested canalizing. $f$ consists of three layers with layer structure $(1,1,2)$, while $g$ possesses only two layers and layer structure $(1,3)$.
@@
A phenotype is associated with a group of attractors where a subset of the system’s variables have a shared state. The states of the shared attractors will be called markers of the phenotype.
@@
Suppose $F=F_{1}\rtimes_{P}F_{2}$ is a decomposable network. Suppose further that only one node $x\in F_{2}$ with update function $f_{x}$ is regulated by nodes in $F_{1}$. If $f_{x}$ is canalizing with $r$ layers, let $\ell\in\{1,\ldots,r\}$ be the lowest (i.e., most important) layer of $f_{x}$, which contains nodes from $F_{1}$. If all regulators of $x$ from $F_{1}$ appear in the core polynomial, we set $\ell=r+1$. Then, setting $y\not\in F_{1}$ to its canalizing value decouples the systems $F_{1}$ and $F_{2}$, as long as $y$ appears in a layer $<\ell$.
@@
The lemma is a direct consequence of Theorem 4.4. If $y$ receives its canalizing input and is in a more important layer of $f_{x}$ than all variables in $F_{1}$, then none of these variables can affect $f_{x}$ anymore. Thus, controlling $y$ to receive its canalizing input eliminates the link between $F_{1}$ and $F_{2}$.
@@
Suppose $F=F_{1}\rtimes_{P_{1}}F_{2}\rtimes_{P_{2}}\cdots\rtimes_{P_{n-1}}F_{n}$ is a decomposable network. If for some $i<j$,
@@
    $f_{x}$ is a canalizing function, which possesses none of the variables from $F_{i}$ in its most important layer, and
@@
    the phenotype of interest depends only on variables in $F_{j}$ and modules that are not “downstream” of $F_{i}$ in the directed acyclic graph of $F$ (see Defintion 2.12),
then the module $F_{i}$ can be excluded from the control search by setting any node $y\notin F_{i}$ to its canalizing input, as long as this node appears in a more dominant layer of $f_{x}$ than all variables of $F_{i}$.
@@
By Lemma 4.7, setting $y$ to its canalizing value results in decoupling $F_{i}$ and $F_{j}$. $F_{i}$ will no longer have any effect on $F_{j}$, and thus, due to condition (iii), on the phenotype of interest. $F_{i}$ can therefore be removed from the control search.
Theorem 4.8 is illustrated in Fig. 4. Note that node $y$ can be in $F_{j}$ or some other module as in the figure.
@@
The method in Theorem 4.8 can be extended to the case when $F_{i}$ and $F_{j}$ are connected via multiple nodes. In that case decoupling is achieved through the same procedure presented above, applied to each node in $F_{j}$ that is regulated by nodes in $F_{i}$.
In Theorem 4.8, we assumed that none of the variables of $F_{i}$ are in the most dominant layer in the update rules of variables in $F_{j}$. If some of the variables of $F_{i}$ are in the most dominant layer, we can still remove module $F_{i}$ from the control search using an edge control, as shown in the following theorem.
@@
Suppose $F=F_{1}\rtimes_{P_{1}}F_{2}\rtimes_{P_{2}}\cdots\rtimes_{P_{n-1}}F_{n}$ is a decomposable network. If for some $i<j$,
@@
    $f_{x}$ is a canalizing function with some variables from $F_{i}$ in its most important layer, and
@@
    the phenotype of interest depends only on variables in $F_{j}$ and modules that are not “downstream” of $F_{i}$ in the directed acyclic graph of $F$ (see Defintion 2.12),
@@
Let $y\in F_{i}$ such that $y\in\mbox{\rm supp}\,(f_{x})$, and that $y$ is located in the most dominant layer $f_{x}$. Then, setting $y$ to its canalizing value results in decoupling the subnetworks $F_{i}$ and $F_{j}$. Thus, $F_{i}$ will no longer have any effect on $F_{j}$ and thus it can be removed from the control search.
@@
The method can be extended to the case when $F_{i}$ and $F_{j}$ are connected via multiple nodes. In that case decoupling is achieved through the same procedure presented above applied to each node in $F_{j}$ with regulators from $F_{i}$.

### v3

@@ 3 Control via Modularity
Therefore,
$\mathcal{A}(F^{\mu})=\mathcal{A}(F_{1}^{\mu_{1}}\rtimes_{P}F_{2}^{\mu_{2}})=\bigsqcup_{\mathcal{C}^{\prime}\in\mathcal{A}(F_{1}^{\mu_{1}})}\mathcal{C}^{\prime}\oplus\mathcal{A}(F^{\mathcal{C}^{\prime},\mu_{2}}_{2})=\mathcal{C}_{1}\oplus\mathcal{A}(F^{\mathcal{C}_{1},\mu_{2}}_{2})=\mathcal{C}_{1}\oplus\mathcal{C}_{2}.$
For the last equality we used the fact that the product of a steady state and a cycle (or vice versa) will result in only one attractor for the combined network. Notice, however, that multiplying two attractors (of length greater than 1) might result in several attractors for the composed network due to the attractors starting at different states.
It follows that there is only one attractor of $F^{\mu}$ and that attractor is $\mathcal{C}_{1}\oplus\mathcal{C}_{2}$. Thus, $F$ is stabilized by $\mu=(\mu_{1},\mu_{2})$ and we have $\mathcal{A}(F^{\mu})=\mathcal{C}$.
Example 3.6.
Consider the network $F(x_{1},x_{2},x_{3},x_{4})=(x_{2},x_{1},x_{2}\wedge x_{4},x_{3})$ whose wiring diagram and state space are given in Figure 4, which can be decomposed into $F=F_{1}\rtimes_{P}F_{2}$, with $F_{1}=F|_{S_{1}}(x_{1},x_{2})=(x_{2},x_{1})$ where $S_{1}=\{x_{1},x_{2}\}$ and $F_{2}=F|_{S_{2}}(x_{3},x_{4})=(e_{2}\wedge x_{4},x_{3})$ where $S_{2}=\{x_{3},x_{4}\}$ and $e_{2}$ is an external parameter. Here the coupling is given by $P=\{x_{2}\to e_{2}\}$. Suppose we want to stabilize $F$ in 0110, which is not an attractor of $F$ (see Figure 4(c)). The set of attractors of $F_{1}$ is given by $\mathcal{A}(F_{1})=\left\{00,11,\{01,10\}\right\}$.
[…]
  • Finally, the control $\mu=(\mu_{1},\mu_{2})$ stabilizes $F$ at $\mathcal{C}=\mathcal{C}_{1}\oplus\mathcal{C}_{2}=\{0110\}$ by Theorem 3.5.
@@
Given a decomposable network $F=F_{1}\rtimes_{P}F_{2}$, let $\mathcal{C}=\mathcal{C}_{1}\oplus\mathcal{C}_{2}$ be an attractor of $F$, where $\mathcal{C}_{1}\in\mathcal{A}(F_{1})$ and $\mathcal{C}_{2}\in\mathcal{A}(F^{\mathcal{C}_{1}}_{2})$ and at least $\mathcal{C}_{1}$ or $\mathcal{C}_{2}$ is a steady state. If $\mu_{1}$ is a control that stabilizes $F_{1}$ in $\mathcal{C}_{1}$ and $\mu_{2}$ is a control that stabilizes $F^{\mathcal{C}_{1}}_{2}$ in $\mathcal{C}_{2}$, then $\mu=(\mu_{1},\mu_{2})$ is a control that stabilizes $F$ in $\mathcal{C}$.
Theorem 3.5 uses the modular structure of a Boolean network to identify controls that stabilize the network in any desired attractor. In biological applications, the attractors typically correspond to distinct biological phenotypes (defined more rigorously in the next section) and a common question is how to force a network to always transition to only one of these phenotypes. For example, cancer biologists may use an appropriate Boolean network model with the two phenotypes proliferation and apoptosis to identify drug targets (i.e., edge or node controls), which force the system to always undergo apoptosis. In [24], we applied the approach described in Corollary 3.7 to find an efficient intervention that solves the control problem for the pancreatic cancer model with 69 nodes. In the following example, we illustrate the approach described in Corollary 3.7 using a toy model.
Example 3.8.
Consider again the network $F(x_{1},x_{2},x_{3},x_{4})=(x_{2},x_{1},x_{2}\wedge x_{4},x_{3})=F_{1}\rtimes_{P}F_{2}$ from Example 3.6 with $F_{1}=F|_{S_{1}}(x_{1},x_{2})=(x_{2},x_{1})$ where $S_{1}=\{x_{1},x_{2}\}$ and $F_{2}=F|_{S_{2}}(x_{3},x_{4})=(e_{2}\wedge x_{4},x_{3})$ where $S_{2}=\{x_{3},x_{4}\}$ and $e_{2}$ is an external parameter. Here the coupling is given by $P=\{x_{2}\to e_{2}\}$. Suppose we want to stabilize $F$ in $1$, which is one of the attractors of $F$ (see Figure 4(c)). The set of attractors of $F_{1}$ is $\mathcal{A}(F_{1})=\left\{00,11,\{01,10\}\right\}$. Let $\mathcal{C}_{1}=\{11\}\in\mathcal{A}(F_{1})$.
[…]
@@
In Theorem 3.5, we required one of the stabilized attractors to be a steady state in order to be able to combine the controls from the modules. We can remove this requirement by using the following definition of stabilization for non-autonomous networks (defined in [24]), which will guarantee that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ can be combined in a unique way, resulting in a unique attractor of the whole network.
@@
[24] A non-autonomous Boolean network is defined by
@@
Consider a controlled non-autonomous network given by $y(t+1)=F_{2}(g(t),y(t),\mu)$, where $g(t)$ is a trajectory representation of an attractor $\mathcal{C}_{1}$ of an upstream network. We say that a control $\mu_{2}$ stabilizes this network, $F_{2}^{\mathcal{C}_{1}}$ (as in Definition 3.10), at an attractor $\mathcal{C}_{2}$ when the resulting network after applying $\mu_{2}$ (denoted here as $F_{2}^{\mathcal{C}_{1},\mu_{2}}$) has $\mathcal{C}_{2}$ as its unique attractor. For non-autonomous networks the definition of unique attractor requires that $\left(g(t),y(t)\right)_{t=0}^{\infty}$ has a unique periodic trajectory up to shifting of $t$ (which is automatically satisfied if $\mathcal{C}_{1}$ or $\mathcal{C}_{2}$ is a steady state).
Figure 4: Wiring diagram, directed acyclic graph, and state space for the network in Examples 3.6 and 3.8. (a) Wiring diagram where the modules are highlighted by amber and green boxes. (b) Directed acyclic graph describing the corresponding connections between the modules. (c) State space of the network generated with Cyclone [40].
@@ 4 Control via Modularity and Canalization
In addition to using the modular structure of the network, we can take advantage of the canalizing structure of the regulatory functions to identify control targets.
@@
$f(\operatorname{\mathbf{x}})\neq f(\operatorname{\mathbf{x}}+\,\mathbf{e_{i}}),$
where $\mathbf{e_{i}}$ is the $i$th unit vector. In that case, we also say $f$ depends on $x_{i}$.
@@
$f(x_{1},x_{2},\ldots,x_{n})=\begin{cases}b,&\ \text{if}\ x_{i}=a\\ g(x_{1},x_{2},\ldots,x_{i-1},x_{i+1},\ldots,x_{n}),&\ \text{if}\ x_{i}\neq a\end{cases}$
@@
A Boolean function $f(x_{1},\ldots,x_{n})$ is a nested canalizing function (NCF) with respect to the permutation $\sigma\in\mathcal{S}_{n}$, inputs $a_{1},\ldots,a_{n}$ and outputs $b_{1},\ldots,b_{n}$, if
$f(x_{1},\ldots,x_{n})=\left\{\begin{array}[c]{ll}b_{1}&x_{\sigma(1)}=a_{1},\\ b_{2}&x_{\sigma(1)}\neq a_{1},x_{\sigma(2)}=a_{2},\\ b_{3}&x_{\sigma(1)}\neq a_{1},x_{\sigma(2)}\neq a_{2},x_{\sigma(3)}=a_{3},\\ \vdots&\vdots\\ b_{n}&x_{\sigma(1)}\neq a_{1},\ldots,x_{\sigma(n-1)}\neq a_{n-1},x_{\sigma(n)}=a_{n},\\ 1+\,b_{n}&x_{\sigma(1)}\neq a_{1},\ldots,x_{\sigma(n-1)}\neq a_{n-1},x_{\sigma(n)}\neq a_{n}.\end{array}\right.$
@@
To account for partial canalization, we also define $k$-canalizing functions, which were first introduced in [41].
Definition 4.4.
A Boolean function $f(x_{1},\ldots,x_{n})$ is $k$-canalizing, where $1\leq k\leq n$, with respect to the permutation $\sigma\in\mathcal{S}_{n}$, inputs $a_{1},\ldots,a_{k}$, and outputs $b_{1},\ldots,b_{k}$ if
$\begin{array}[]{l}f(x_{1},\ldots,x_{n})=\\ \left\{\begin{array}[c]{ll}b_{1}&x_{\sigma(1)}=a_{1},\\ b_{2}&x_{\sigma(1)}\neq a_{1},x_{\sigma(2)}=a_{2},\\ b_{3}&x_{\sigma(1)}\neq a_{1},x_{\sigma(2)}\neq a_{2},x_{\sigma(3)}=a_{3},\\ \vdots&\vdots\\ b_{k}&x_{\sigma(1)}\neq a_{1},\ldots,x_{\sigma(k-1)}\neq a_{k-1},x_{\sigma(k)}=a_{k},\\ f_{C}\not\equiv b_{k}&x_{\sigma(1)}\neq a_{1},\ldots,x_{\sigma(k-1)}\neq a_{k-1},x_{\sigma(k)}\neq a_{k},\end{array}\right.\end{array}$
where $f_{C}=f_{C}(x_{\sigma(k+1)},\ldots,x_{\sigma(n)})$ is the core function, a Boolean function on $n-k$ variables. When $f_{C}$ is not canalizing, then the integer $k$ is the canalizing depth of $f$ [42]. Note that an $n$-canalizing function (i.e., a function where all variables become eventually canalizing) is an NCF.
@@
where each $M_{i}=\displaystyle\prod_{j=1}^{k_{i}}(x_{i_{j}}+a_{i_{j}})$ is a nonconstant function, called an extended monomial, $p_{C}$ is the core polynomial of $f$, and $k=\displaystyle\sum_{i=1}^{r}k_{i}$ is the canalizing depth. Each $x_{i}$ appears in exactly one of $\{M_{1},\ldots,M_{r},p_{C}\}$, and the only restrictions are the following “exceptional cases”:
@@
Given a Boolean function $f(x_{1},\ldots,x_{n})$ represented as in Equation 3, we call the extended monomials $M_{i}$ the layers of $f$ and if $i<j$, we say that $M_{i}$ and its variables are more dominant and $M_{j}$ and its variables are less dominant. We also define, as in [43], the layer structure as the vector $(k_{1},\ldots,k_{r})$, which describes the number of variables in each layer.
Note that $f$ is nested canalizing if and only if $k_{1}+\cdots+k_{r}=n$.
Remark 4.7.
Here we note the following important properties of layers of canalization.
  1. (a)
    Theorem 4.5 shows that any Boolean function has a unique extended monomial form given by Equation 3, in which the variables are partitioned into different layers based on their dominance. Any variable that is canalizing (independent of the values of other variables) is in the first layer. Any variable that “becomes” canalizing when excluding all variables from the first layer is in the second layer, etc. All remaining variables that never become canalizing are part of the core polynomial. The number of variables that eventually become canalizing is the canalizing depth of the function. NCFs are exactly those functions where all variables become eventually canalizing (note not all variables of an NCF must be in the first layer).
  2. (b)
    While variables in the same layer may have different canalizing input values, they all share the same canalized output value, i.e., they all canalize a function to the same output. On the other hand, the outputs of two consecutive layers are distinct. Therefore, the number of layers of a $k$-canalizing function, expressed as in Definition 4.4, is simply one plus the number of changes in the vector of canalized outputs, $(b_{1},b_{2},\ldots,b_{k})$.
@@
The Boolean functions
$f(x_{1},x_{2},x_{3},x_{4})=x_{1}\wedge(\neg x_{2}\vee(x_{3}\wedge x_{4}))\qquad\text{and}\qquad g(x_{1},x_{2},x_{3},x_{4})=x_{1}\wedge(\neg x_{2}\vee x_{3}\vee x_{4})$
are nested canalizing. The function $f$ has layer structure $(1,1,2)$ because
  • $x_{1}$ canalizes $f$ to 00 if it receives its canalizing input 00;
  • if this is not the case, $x_{2}=0$ canalizes $f$ to $1$;
  • $x_{3}=0$ or $x_{4}=0$ canalizes $f$ to 00.
On the other hand, $g$ has layer structure $(1,3)$. As for $f$, $x_{1}=0$ canalizes $g$ to 00. If this does not happen, any of the following, $x_{2}=0,x_{3}=1,x_{4}=1$, canalizes $g$ to $1$.
@@
Next we will define phenotypes. In meaningful biological networks, the attractors correspond to phenotypes. Typically a small subset of all Boolean variables is used to define phenotypes.
Definition 4.9.
Given a Boolean network $F$ with attractor space $\mathcal{A}(F)$ and phenotype-defining variables $\mathcal{P}\subset\{x_{1},\ldots,x_{n}\}$, we associate the same phenotype to all attractors $\mathcal{C}\in\mathcal{A}(F)$ that are identical in $\mathcal{P}$. The states in $\mathcal{P}$ will be called markers of the phenotype.
@@
By Theorem 4.5, the variables of any Boolean update function can be ordered by importance/dominance, based on which layer they appear in, which is why in Definition 4.6 we called variables in lower/upper layers “more/less dominant,” respectively. Thus, once we control a variable in a certain layer (by setting it to its canalizing input value), any further control of variables in less dominant layers will have no effect on the function (and thus on the network). We state this fact in the following lemma.
@@
Suppose $F=F_{1}\rtimes_{P}F_{2}$ is a decomposable network. Suppose further that only one node $x\in F_{2}$ with update function $f_{x}$ is regulated by nodes in $F_{1}$. If $f_{x}$ is canalizing with $r$ layers, let $M_{\ell}$ be the most dominant layer of $f_{x}$ which contains nodes from $F_{1}$. If all regulators of $x$ from $F_{1}$ appear in the core polynomial, we set $\ell=r+1$. Then, setting $y\not\in F_{1}$ to its canalizing value decouples the systems $F_{1}$ and $F_{2}$, as long as $y$ appears in a layer of $f_{x}$ that is more dominant than $M_{\ell}$.
@@
The lemma is a direct consequence of Theorem 4.5. If $y$ receives its canalizing input and is in a more dominant layer of $f_{x}$ than all variables in $F_{1}$, then none of these variables can affect $f_{x}$ anymore. Thus, controlling $y$ to receive its canalizing input eliminates the link between $F_{1}$ and $F_{2}$.
The modularity approach in [24] yields an acyclic directed graph after one collapses each module into a single node. This endows a natural partial ordering on the collection modules of a network where one module precedes the next if there is a path for every node in the first module which ends in the second module. While not all modules are comparable, we can speak of chains of modules which consist of subsets of the partial ordering which are totally ordered. Furthermore, one can rank the modules based on the percentile scores (i.e., rank module $k$ out of $m$ modules). This type of ranking has been studied in [45], where it was shown that the importance of the modules is strongly correlated with the aggressiveness of mutations occurring within those modules and the effectiveness of interventions.
@@
Suppose $F=F_{1}\rtimes_{P_{1}}F_{2}\rtimes_{P_{2}}\cdots\rtimes_{P_{n-1}}F_{m}$ is a decomposable network. If for some $i<j$,
@@
    $f_{x}$ is a canalizing function, which possesses none of the variables from $F_{i}$ in its most dominant layer, and
@@
    the phenotype of interest only depends on variables in $F_{j}$ as well as modules $F_{k}$, for which any chain containing $F_{i}$ and $F_{k}$ also contains $F_{j}.$
then the module $F_{i}$ can be excluded from the control search by setting any node $y\notin F_{i}$ to its canalizing input, as long as this node appears in a more dominant layer of $f_{x}$ than all inputs of $f_{x}$ that are part of $F_{i}$.
@@
By Lemma 4.10, setting $y$ to its canalizing value results in decoupling $F_{i}$ and $F_{j}$. $F_{i}$ will no longer have any effect on $F_{j}$, and thus, due to condition (iii), on the phenotype of interest. Therefore, $F_{i}$ can be removed from the control search.
Figure 5: Example of a modular directed acylic graph structure to illustrate Condition (iii) in Theorems 4.11 and 4.13. (a) Module $F_{i}$ can be removed from the control search as long as conditions (i) and (ii) are satisfied, and the phenotype of interest depends only on any subset of variables that are part of the blue modules. (b) When a node in module $F_{m}$ is regulated by a node in module $F_{k}$ (indicated by the red edge in the directed acyclic graph), the phenotype may no longer depend on nodes in $F_{m}$, in order for module $F_{i}$ to be removable from the control search.
The third condition of this theorem is illustrated in Figure 5.
@@
The method in Theorem 4.11 can be extended to the case when $F_{i}$ and $F_{j}$ are connected via multiple nodes. In that case, decoupling is achieved through the same procedure presented above, applied to each node in $F_{j}$ that is regulated by nodes in $F_{i}$.
Figure 6: Control via modularity and canalization. Once the network is decomposed into modules $F_{1},\ldots,F_{m}$, we can override the effect of module $F_{i}$ by the using another module ($F_{k}$ in this case) whose variables are inputs of $f_{x}$ that are located in more dominant layers than the layers containing the variables of $F_{i}$.
Theorem 4.11 is illustrated in Figure 6. Note that node $y$ can be in $F_{j}$ or some other module as in the figure. In this theorem, we assumed that none of the variables of $F_{i}$ are in the most dominant layer in the update rules of variables in $F_{j}$. If some variables of $F_{i}$ are in the most dominant layer, we can still remove module $F_{i}$ from the control search using an edge control, as shown in the following theorem.
@@
Suppose $F=F_{1}\rtimes_{P_{1}}F_{2}\rtimes_{P_{2}}\cdots\rtimes_{P_{n-1}}F_{m}$ is a decomposable network. If for some $i<j$,
@@
    $f_{x}$ is a canalizing function with some variables from $F_{i}$ in its most dominant layer, and
@@
    the phenotype of interest only depends on variables in $F_{j}$ as well as modules $F_{k}$, for which any chain containing $F_{i}$ and $F_{k}$ also contains $F_{j}.$
@@
Let $y\in F_{i}$ such that $y$ appears as a variable in $f_{x}$, and that $y$ is located in the most dominant layer $f_{x}$. Then, setting $y$ to its canalizing value results in decoupling the subnetworks $F_{i}$ and $F_{j}$. Thus, $F_{i}$ will no longer have any effect on $F_{j}$ and thus it can be removed from the control search.
@@
The method can be extended to the case when $F_{i}$ and $F_{j}$ are connected via multiple nodes (that is, condition (i) in Theorem 4.11 and Theorem 4.13 can be relaxed. In that case decoupling is achieved through the same procedure presented above, applied to each node in $F_{j}$ with regulators from $F_{i}$.
We further note that condition (ii) in the theorems above is generally very restrictive as only a small proportion of Boolean functions in $n>3$ variables are canalizing, let alone nested canalizing. However, as shown in [30], most biological Boolean network models are almost entirely governed by nested canalizing functions.
5 An application: Control of a Blood Cancer Boolean Model